# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 8 Densification — Phase 1
Let $\alpha=\frac{1}{3}-2s>0$, then we show that $\rho=\alpha\varepsilon$ suffices. Set $C:=5$, $s^{\prime}=\frac{s}{1-\rho}$, then $s^{\prime}<1/6$ and $6s^{\prime}<1-2\alpha$. Let $c_{0}=c_{0}(s^{\prime},8C)$ be the constant guaranteed by Lemma 8.1. Then there exists some $c_{1}=c_{1}(s^{\prime},\varepsilon)>0$ such that for every $q\in(0,c_{1})$, we have $c_{0}q^{6s^{\prime}}>q^{1-2\alpha}$.
We define the sequence of progressively denser induced subgraphs $G=G_{0}\supset G_{1}\supset\dots$ as follows. Assume that $G_{k}$ is already defined on $n_{k}$ vertices with edge density $p_{k}$. If $p_{k}>c_{1}$ or $n_{k}<n^{1-\varepsilon}$, we stop, otherwise we define $G_{k+1}$ as follows. By Lemma 8.2 (noting that the condition $\operatorname{surp}(G)\leq|\lambda_{n}|n<0.01pn^{2}$ is trivially satisfied), either (i) $G_{k}$ contains a subgraph $G^{\prime}$ on at least $n_{k}/2$ vertices with average degree at least $p_{k}n_{k}/4$ and $\Delta(G^{\prime})\leq Cp_{k}n_{k}$, or (ii) $G$ contains a subgraph $G^{\prime\prime}$ on at least $n_{k}/C$ vertices of average degree at least $0.2Cp_{k}n_{k}$.
First, consider the case if outcome (ii) happens. Then set $G_{k+1}:=G^{\prime\prime}$, and note the $n_{k+1}\geq n_{k}/C$ and $p_{k+1}\geq 0.2C^{2}p_{k}>Cp_{k}$.
Now consider the case if outcome (i) happens. Let $n^{\prime}=v(G^{\prime})$ and let $p^{\prime}$ be the edge density of $G^{\prime}$, then $n^{\prime}\geq n_{k}/2$, $p^{\prime}\geq p_{k}/4$ and $\Delta(G^{\prime})\leq Cp_{k}n_{k}\leq 8Cp^{\prime}n^{\prime}$. As $n_{k}\geq n^{1-\varepsilon}$, we have $p^{\prime}\geq p/4\geq n^{-\rho}/4\geq(n^{\prime})^{-\rho/(1-\varepsilon)}/8>(n^{\prime})^{-1/2}$. Also, if $\lambda^{\prime}$ is the smallest eigenvalue of $G^{\prime}$, then
$|\lambda^{\prime}|\leq|\lambda_{n}|\leq n^{s}\leq(n^{\prime}/2)^{s/(1-\rho)}\leq(n^{\prime})^{s^{\prime}}.$
Hence, we can apply Lemma 8.1 to conclude that $G^{\prime}$ contains a subgraph on at least $p^{\prime}n^{\prime}$ vertices of edge density at least $c_{0}(p^{\prime})^{6s^{\prime}}$. Let $G_{k+1}$ be this subgraph. Then $n_{k+1}\geq p^{\prime}n^{\prime}\geq p_{k}n_{k}/8$ and
$p_{k+1}\geq c_{0}(p^{\prime})^{6s^{\prime}}\geq(p_{k}/8)^{1-2\alpha}.$
Here, the last inequality holds by our assumption that $p_{k}\leq c_{1}$.
We can unite outcomes (i) and (ii) by observing that in both cases there is some $q_{k}>1$ such that $n_{k+1}>n_{k}/q_{k}$ and $p_{k+1}>p_{k}q_{k}^{2\alpha}$. Indeed, in case (i), we can choose $q_{k}=C$, and in case (ii), we choose $q_{k}=p_{k}$. From this, it is easy to see that the sequence $n_{k}^{2\alpha}p_{k}$ is monotone increasing, so we have $n_{k}^{2\alpha}p_{k}\geq n^{2\alpha}p$. Hence, as long as $p_{k}\leq c_{1}$, we have $n_{k}\geq n(p/c_{1})^{1/2\alpha}>n^{1-\rho/2\alpha}/c_{1}^{1/2\alpha}=n^{1-\varepsilon/2}/c_{1}^{1/2\alpha}>n^{1-\varepsilon}$. Let $K$ be the last index $k$ for which $G_{k}$ is defined, so either $p_{K}>c_{1}$ or $n_{K}<n^{1-\varepsilon}$. Then the previous argument shows that we must have $p_{K}>c_{1}$ and $n_{K}>n^{1-\varepsilon}$, so the graph $G_{K}$ suffices. ∎
@@
Next, we prove a version of the previous lemmas for small surplus. As the proofs are more or less the same, with only some parameters changed, we only highlight the key differences.
@@
Let $s\in(0,1/60)$, $C>2$, and let $G$ be a $n$-vertex graph with edge density $p>n^{-1/3}$, $\Delta(G)\leq Cpn$, and $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Then $G$ has a subgraph on at least $pn$ vertices of edge density at least $c_{0}p^{4/5}$ for some $c_{0}=c_{0}(s,C)>0$.
@@
By Lemma 5.3, (i), we have
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{0<\lambda_{i}}|\lambda_{i}|\leq\operatorname{surp}^{*}(G)\leq n^{1+s}.$
By Lemma 6.5, we also have
@@
for every $T>C_{0}n^{1-\frac{1}{24}+\frac{s}{4}}$. So we can apply Lemma 6.7 with $q=3/80>s$ and the sequence of positive eigenvalues to get that for every $H\leq n$,
$\sum_{0<\lambda_{i}\leq H}\lambda_{i}\leq cn^{1+s/q}H^{1-s/q}.$
Furthermore, by Lemma 5.3, (ii), we have
$\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq O(n^{1/2}\operatorname{surp}^{*}(G))=O(n^{3/2+s}).$
Write $u:=s/q<4/9$, then setting $H:=(8c)^{-1/(1-u)}p^{1/(1-u)}n$, we get $\sum_{\lambda_{i}<H}\lambda_{i}^{2}\leq pn^{2}/4$. Thus, if $N$ is the number of triangles, then
$6N=\sum_{i=1}^{n}\lambda_{i}^{3}\geq H\sum_{\lambda_{i}\geq H}\lambda_{i}^{2}-\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\geq Hpn^{2}/2-O(n\operatorname{surp}^{*}(G))\geq c^{\prime}p^{\frac{2-u}{1-u}}n^{3}-n^{2+s}.$
Here, in the second inequality, we used Lemma 5.3, (iii). The rest of the proof is identical to the proof of Lemma 8.1. ∎
@@
Let $s\in(0,1/60)$, $\varepsilon\in(0,1/2)$, then there exists $\rho=\rho(s,\varepsilon)>0$ and $c_{1}=c_{1}(s,\varepsilon)>0$ such that the following holds for every $n>n_{0}(s,\varepsilon)$. Let $G$ be an $n$-vertex graph with edge density $p$ such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$ and $p>n^{-\rho}$. Then $G$ has a subgraph on at least $n^{1-\varepsilon}$ vertices with edge density at least $c_{1}$.
@@
The proof of this is almost identical to the proof of Lemma 8.3, but we use Lemma 8.4 instead of Lemma 8.1. We omit further details. ∎
9 Densification — Phase 2
In this section, we prove that graphs of positive constant density and large smallest eigenvalue (or small surplus) are $o(1)$-close to the disjoint union of cliques. In particular, this implies that such graphs must contain subgraphs of density 1o(1)1-o(1). Beyond the use of the main lemmas, Lemma 6.3 and 6.5, another key component is showing that if the adjacency matrix of a graph is close to a small rank matrix (in Frobenius norm), then $G$ admits an ultra-strong regularity partition. These kind of partitions are closely related to Szemerédi’s regularity lemma, but they provide substantially stronger quantitative bounds. Ultra-strong regularity lemmas first appeared in relation to graphs of bounded VC-dimension, see the seminal work of Lovász and Szegedy [60].
A partition $V_{1},\dots,V_{K}$ of a set of size $n$ is an equipartiton, if $|V_{i}|\in\{\lfloor n/K\rfloor,\lceil n/K\rceil\}$ for $i\in[K]$. Given a graph $G$, $\delta\in(0,1)$, and two disjoint sets $X,Y\subset V(G)$, the pair $(X,Y)$ is $\delta$-empty if there are at most $\delta|X||Y|$ edges between $X$ and $Y$. Also, $(X,Y)$ is $\delta$-full if there are at least $(1-\delta)|X||Y|$ edges between $X$ and $Y$. Then $(X,Y)$ is $\delta$-homogeneous if it is either $\delta$-empty or $\delta$-full.
A $\delta$-regular partition of $G$ is an equipartition $V_{1},\dots,V_{K}$ of the vertex set such that all but at most $\delta K^{2}$ of the pairs $(V_{i},V_{j})$ for $1\leq i<j\leq K$ are $\delta$-homogeneous.
@@ 9 Densification — Phase 2
For every $\delta>0$, there exists $\varepsilon>0$ such that the following holds for every positive integer $r$, and every $n$ that is sufficiently large with respect to $\delta,r$. Let $G$ be a graph with adjacency matrix $A$. Assume that there exists an $n\times n$ symmetric matrix $B$ of rank $r$ such that $A-B||_{F}^{2}\leq\varepsilon n^{2}$. Then $G$ has a $\delta$-regular partition into $K$ parts, where $1/\delta<K<O_{r,\delta}(1)$.
@@
Let $\beta:=10^{-3}\delta^{1/2}r^{-3/2}$, and for $\ell\in\mathbb{Z}$, let
@@
$\bigcup_{\overline{\ell}\in I}X_{\overline{\ell}}\geq n\left(1-\frac{\delta}{8r}\right)^{r}\geq n(1-\delta/8).$ (6)
@@
$\gamma=\gamma_{\overline{\ell}_{1},\overline{\ell}_{2}}=\sum_{i=1}^{r}\mu_{i}\cdot\frac{\beta^{2}}{n}\cdot\overline{\ell}_{1}(i)\overline{\ell}_{2}(i).$
@@
$\left|w_{i}(j_{1})w_{i}(j_{2})-\frac{\beta^{2}}{n}\overline{\ell_{1}}(i)\overline{\ell_{2}}(i)\right|\leq 4h,$
which we get from the general inequality $|ab-cd|\leq|a||b-d|+|d||a-c|$. Therefore, we have
$|B_{j_{1},j_{2}}-\gamma|\leq\sum_{i=1}^{r}|\mu_{i}|\cdot\left|w_{i}(j_{1})w_{i}(j_{2})-\frac{\beta^{2}}{n}\overline{\ell_{1}}(i)\overline{\ell_{2}}(i)\right|\leq\sum_{i=1}^{r}|\mu_{i}|\cdot\frac{\beta^{2}}{n}\cdot 4h\leq 8r\beta^{2}h<\frac{1}{3}.$ (7)
Observe that if $X\subset X_{\overline{\ell}_{1}}$ and $Y\subset X_{\overline{\ell}_{2}}$ are such that $(X,Y)$ is not $\delta$-homogeneous, then
$A[X\times Y]-B[X\times Y]||_{F}^{2}\geq\frac{\delta}{36}|X||Y|.$
Indeed, if $\gamma_{\overline{\ell}_{1},\overline{\ell}_{2}}<1/2$, then $A_{j_{1},j_{2}}-B_{j_{1},j_{2}}\geq 1/6$ for every one entry $A_{j_{1},j_{2}}$, otherwise $|A_{j_{1},j_{2}}-B_{j_{1},j_{2}}|\geq 1/6$ for every zero entry $A_{j_{1},j_{2}}$.
Now let $K=|I|/(8\delta)$, and let $V_{1},\dots,V_{K}$ be an equipartition of $V(G)$ as follows. Let $V^{*}$ be the set of elements not covered by any of the $X_{\overline{\ell}}$ for $\overline{\ell}\in I$. For each $\overline{\ell}\in I$, choose as many of the $V_{i}$ to be completely contained in $X_{\overline{\ell}}$ as possible, and then move the not covered elements of $X_{\overline{\ell}}$ to $V^{*}$. Then finally partition $V^{*}$. Each $X_{\overline{\ell}}$ contributes at most $n/K$ elements to $V^{*}$, so in the end we have $|V^{*}|\leq\delta n/8+|I|\cdot(n/K)\leq\delta n/4$. Therefore, at most $\delta K/4$ sets $V_{i}$ are contained in $V^{*}$. We show that $V_{1},\dots,V_{K}$ is a $\delta$-regular partition.
Assume that $(V_{i},V_{j})$ is not $\delta$-homogeneous. There are at most $\delta K^{2}/2$ such pairs where either $V_{i}\subset V^{*}$ or $V_{j}\subset V^{*}$. On the other hand, if $V_{i},V_{j}\not\in V^{*}$, then $A[V_{i}\times V_{j}]-B[V_{i}\times V_{j}]||_{F}^{2}\geq\frac{\delta}{36}|V_{i}||V_{j}|$. As $A-B||_{F}^{2}\leq\varepsilon n^{2}$, this means that the number of such pairs is at most $36\varepsilon/\delta K^{2}\leq\delta K^{2}/2$. Hence, the total number of pairs that are not $\delta$-homogeneous is at most $\delta K^{2}$, as desired. ∎
An important feature of Lemma 9.1 that $\varepsilon$ only depends on $\delta$, and not on $r$. To continue from this point, we observe that if $X,Y,Z$ are sets of size $\Omega(n)$ such $(X,Y)$ and $(Y,Z)$ are $\delta$-full, then $(X,Z)$ cannot be $\delta$-empty, assuming $\operatorname{surp}^{*}(G)$ is small.

### v2

@@ 7 Densification — Phase 1
We may and will assume that $n$ is sufficiently large. For a graph $H$, we write $v(H)$ for the number of its vertices, $d(H)$ for its average degree, and $p(H)$ for its density.
Let $\gamma^{\prime}\in(\gamma,\frac{1-\varepsilon}{6})$ be any constant such that $\frac{\rho}{\varepsilon}+\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}<1$; then $|\lambda_{n}|\leq n^{\gamma^{\prime}}/2$. Let $H$ be an induced subgraph of $G$ that maximizes the function
$v(H)^{\rho/\varepsilon}\cdot p(H).$
We show that $H$ is the desired subgraph. Due to the maximality, $v(H)^{\rho/\varepsilon}\cdot p(H)\geq v(G)^{\rho/\varepsilon}\cdot p(G)>n^{\rho/\varepsilon-\rho}$. Since $p(H)\leq 1$, we get $v(H)\geq n^{1-\varepsilon}$, as desired. In what follows, we lower bound the density of $H$. Recall that Cauchy’s interlacing theorem ensures that the smallest eigenvalue of $H$ is at least $n^{\gamma^{\prime}}/2$.
First, we show that $p(H)>v(H)^{-\rho}$. Indeed, if $p(H)\leq v(H)^{-\rho}$, then
$v(H)^{\rho/\varepsilon-\rho}\geq v(H)^{\rho/\varepsilon}\cdot p(H)\geq v(G)^{\rho/\varepsilon}\cdot p(G)>n^{\rho/\varepsilon-\rho}.$
This contradicts $v(H)\leq n$. Therefore, $p(H)>v(H)^{-\rho}$. Next, we prove that $d(H)>n^{3\gamma^{\prime}}$. Recall that $\rho<\frac{1}{2},\gamma^{\prime}<\frac{1-\varepsilon}{6}$ and $v(H)>n^{1-\varepsilon}$ is sufficiently large. We have
$d(H)=\big(v(H)-1\big)p(H)>\big(v(H)-1\big)v(H)^{-\rho}>v(H)^{1/2}>n^{(1-\varepsilon)/2}>n^{3\gamma^{\prime}}.$
Now apply Lemma˜7.2 to $H$ with $C=5$. The requirement $\operatorname{surp}(H)\leq\frac{1}{100}d(H)v(H)$ trivially holds as by ˜5.1, we have $\operatorname{surp}(H)\leq n^{\gamma^{\prime}}\cdot v(H)/4<\frac{1}{100}d(H)v(H)$. By Lemma˜7.2, either (i) $H$ contains an induced subgraph $H_{0}$ such that $v(H_{0})=v(H)/5$ and $d(H_{0})\geq d(H)$, or (ii) $H$ contains an induced subgraph $H_{1}$ such that $v(H_{1})\geq v(H)/2,d(H_{1})\geq d(H)/4$ and $\Delta(H_{1})\leq 5d(H)$.
Assume that (i) holds. Using that $v(H)>n^{1-\varepsilon}$ is sufficiently large, we have
$p(H_{0})=\frac{d(H_{0})}{v(H_{0})-1}\geq\frac{d(H)}{v(H)/5-1}=(1-o(1))5p(H).$
Therefore,
$v(H_{0})^{\rho/\varepsilon}\cdot p(H_{0})\geq 5^{-\rho/\varepsilon}v(H)^{\rho/\varepsilon}\cdot(1-o(1))5p(H)=(1-o(1)){5^{1-\rho/\varepsilon}}\cdot v(H)^{\rho/\varepsilon}\cdot p(H).$
As $\rho/\varepsilon<1$, this contradicts the maximality of $H$.
Hence, (ii) must hold. Then, we apply Lemma˜7.1 to $H_{1}$ with $C=20$. To be able to apply this lemma, we first verify the requirements on $p(H_{1})$, $\Delta(H_{1})$ and the smallest eigenvalue. Since $\rho<1/2$ and $v(H_{1})\geq v(H)/2$, we have $p(H)>v(H)^{-\rho}>8\cdot v(H_{1})^{-1/2}$. Thus,
$p(H_{1})=\frac{d(H_{1})}{v(H_{1})-1}\geq\frac{d(H)}{4v(H)}\geq\frac{p(H)}{8}>\frac{1}{8}v(H)^{-\rho}>v(H_{1})^{-1/2}.$
Furthermore, $\Delta(H_{1})\leq 5d(H)\leq 20d(H_{1})$. In addition, the smallest eigenvalue of $H_{1}$ has absolute value at most $n^{\gamma^{\prime}}/2$. As $v(H_{1})\geq v(H)/2>n^{1-\varepsilon}/2$ and $\gamma^{\prime}/(1-\varepsilon)<1$, this is at most $\big(2v(H_{1})\big)^{\gamma^{\prime}/(1-\varepsilon)}/2<v(H_{1})^{\gamma^{\prime}/(1-\varepsilon)}$. Here, $\frac{\gamma^{\prime}}{1-\varepsilon}<1/6$ as $\gamma^{\prime}<\frac{1-\varepsilon}{6}$. Therefore, we can apply Lemma˜7.1 to get an induced subgraph $H_{2}$ of $H_{1}$ such that $v(H_{2})\geq p(H_{1})v(H_{1})$ and $p(H_{2})\geq c_{0}\cdot p(H_{1})^{\frac{2\gamma^{\prime}/(1-\varepsilon)}{1-4\gamma^{\prime}/(1-\varepsilon)}}=c_{0}p(H_{1})^{\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}}$, where $c_{0}=c_{0}(\gamma^{\prime},\varepsilon)>0$. Hence,
$v(H_{2})^{\rho/\varepsilon}\cdot p(H_{2})\geq\big[(p(H_{1})v(H_{1})\big]^{\rho/\varepsilon}\cdot\Big[c_{0}\cdot p(H_{1})^{\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}}\Big]=c_{0}\cdot v(H_{1})^{\rho/\varepsilon}\cdot p(H_{1})^{\frac{\rho}{\varepsilon}+\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}}.$
Recall that $v(H_{1})\geq v(H)/2$ and $p(H_{1})\geq p(H)/5$. Hence,
$v(H_{2})^{\rho/\varepsilon}\cdot p(H_{2})\geq c_{0}^{\prime}\cdot v(H)^{\rho/\varepsilon}\cdot p(H)^{\frac{\rho}{\varepsilon}+\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}}$
for some $c_{0}^{\prime}=c_{0}^{\prime}(\gamma^{\prime},\varepsilon,\rho)>0$. Since $H$ maximizes the function $v(H)^{\rho/\varepsilon}\cdot p(H)$, we have
$v(H)^{\rho/\varepsilon}\cdot p(H)\geq v(H_{2})^{\rho/\varepsilon}\cdot p(H_{2})\geq c_{0}^{\prime}\cdot v(H)^{\rho/\varepsilon}\cdot p(H)^{\frac{\rho}{\varepsilon}+\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}}.$
Using the condition $1>\frac{\rho}{\varepsilon}+\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}$, this implies that $p(H)\geq(c_{0}^{\prime})^{1/\big(1-\frac{\rho}{\varepsilon}-\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}\big)}$. This completes the proof by taking $c_{1}=(c_{0}^{\prime})^{1/\big(1-\frac{\rho}{\varepsilon}-\frac{2\gamma^{\prime}}{1-\varepsilon-4\gamma^{\prime}}\big)}$. ∎
@@
Next, we prove a counterpart of Lemma˜7.3 for the surplus. For this purpose, we require a surplus version of Lemma˜7.1 as follows. As the proofs are more or less the same, with only some parameters changed, we only highlight the key differences.
@@
Let $\gamma\in(0,1/60)$, $C>2$, and let $G$ be a $n$-vertex graph with edge density $p>n^{-1/3}$, $\Delta(G)\leq Cpn$, and $\operatorname{surp}(G)\leq n^{1+\gamma}$. Then $G$ has a subgraph on at least $pn$ vertices of edge density at least $c_{0}p^{4/5}$ for some $c_{0}=c_{0}(\gamma,C)>0$.
@@
Assume $n$ is sufficiently large. By lemma˜5.2, $\operatorname{surp}^{*}(G)=O(\operatorname{surp}(G)\log n)<n^{1+\gamma^{\prime}}$ for some constant $\gamma^{\prime}\in(\gamma,1/60)$. Then, Lemma˜5.3 (i) implies
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{0<\lambda_{i}}|\lambda_{i}|\leq\operatorname{surp}^{*}(G)\leq n^{1+\gamma^{\prime}}.$
By Lemma˜6.6, we also have
@@
for every $T>C_{0}n^{1-\frac{1}{24}+\frac{\gamma^{\prime}}{4}}$. So we can apply Lemma˜6.9 with $q=3/80>\gamma^{\prime}$ and the sequence of positive eigenvalues to get that for every $\kappa\leq 1$,
$\sum_{0<\lambda_{i}\leq\kappa n}\lambda_{i}\leq c\kappa^{1-\gamma^{\prime}/q}n^{2}.$
Furthermore, by Lemma˜5.3 (ii), we have
$\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq O\big(n^{1/2}\operatorname{surp}^{*}(G)\big)=O(n^{3/2+\gamma^{\prime}}).$
Write $u:=\gamma^{\prime}/q<4/9$, and set $\kappa:=(8c)^{-1/(1-u)}p^{1/(1-u)}$. We get $\sum_{\lambda_{i}<\kappa n}\lambda_{i}^{2}\leq pn^{2}/4$. Thus, if $N$ is the number of triangles, then
$6N=\sum_{i=1}^{n}\lambda_{i}^{3}\geq\kappa n\sum_{\lambda_{i}\geq H}\lambda_{i}^{2}-\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}\geq\kappa pn^{3}/2-O\big(n\operatorname{surp}^{*}(G)\big)\geq c^{\prime}p^{\frac{2-u}{1-u}}n^{3}-n^{2+\gamma^{\prime}}.$
Here, we used Lemma˜5.3 (iii) in the second inequality. The rest of the proof is identical to that of Lemma˜7.1. ∎
Now, we state the counterpart of Lemma˜7.3 for surplus – the other main result of this section. The proof of Lemma˜7.5 follows almost identically to that of Lemma˜7.3, except that we use Lemma˜7.4 in place of Lemma˜7.1.
@@
Let $\gamma\in(0,1/60)$, $\varepsilon\in(0,1/2)$. Then there exist $\rho=\rho(\gamma,\varepsilon)>0$ and $c_{1}=c_{1}(\gamma,\varepsilon)>0$ such that the following holds for every $n>n_{0}(\gamma,\varepsilon)$. Let $G$ be an $n$-vertex graph with edge density $p>n^{-\rho}$ and $\operatorname{surp}(G)\leq n^{1+\gamma}$. Then $G$ has a subgraph on at least $n^{1-\varepsilon}$ vertices with edge density at least $c_{1}$.
8 Densification — Phase 2
In this section, we prove that graphs of positive constant density and large smallest eigenvalue (or small surplus) are $o(1)$-close to the disjoint union of cliques. In particular, this implies that such graphs must contain subgraphs of density 1o(1)1-o(1).
In the first step of the proof, we use our main lemmas, Lemma 6.3 and 6.6, to show that the adjacency matrix of a graph is close to a constant-rank matrix, in Frobenius norm.
In the second step, we will use this approximation to show that $G$ admits an ultra-strong regularity partition. Ultra-strong regularity partition is a partition of $V(G)$ where almost all pairs of parts have very high or very low density of edges between them. These are closely related to Szemerédi’s regularity lemma, but they provide substantially stronger quantitative bounds. Ultra-strong regularity lemmas first appeared in relation to graphs of bounded VC-dimension; see the seminal work of Lovász and Szegedy [69]. Our approach to finding this regularity partition is morally similar to the spectral approach of Frieze and Kannan [45] (see also [84]), but the good understanding of the spectrum of $G$ coming from Section 6 allows us to obtain a much stronger quantitative result.
Finally, in the last step of the proof, we analyse the regularity partition obtained from the previous step and show that it contains very few induced paths of length 2, i.e. cherries. This shows that the whole graph is close to the union of cliques. We now give the details.
Lemma 8.1.
Let $\gamma\in(0,1/4)$, $\varepsilon>0$, and let $n$ be sufficiently large. If $G$ is an $n$-vertex graph with adjacency matrix $A$ and with $|\lambda_{n}|\leq n^{\gamma}$, then there is a matrix $B$ of rank $O_{\gamma,\varepsilon}(1)$ such that $\|A-B\|_{F}^{2}\leq\varepsilon n^{2}$.
@@ 8 Densification — Phase 2
Let $A=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}$ be the spectral decomposition of $A$. We have
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq|\lambda_{n}|n\leq n^{1+\gamma}.$
Also, Lemma˜6.3 implies $4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}$ for every $T\geq 2n^{1/2+\gamma}\geq 2|\lambda_{n}|\sqrt{n}$. Hence, we can apply Lemma 6.9 to the sequence of positive eigenvalues with $q:=1/4>\gamma$ to conclude that for every $\kappa\in(0,1)$, we have
$\sum_{0<\lambda_{i}<\kappa n}\lambda_{i}^{2}\leq O(\kappa^{1-4\gamma}n^{2}).$
$\sum_{\lambda_{i}<\kappa n}\lambda_{i}^{2}\leq O(\kappa^{1-4\gamma}n^{2})+n^{3/2}.$
Hence, we can choose $\kappa$ (depending only on $\gamma$ and $\varepsilon$) such that $\sum_{\lambda_{i}<\kappa n}\lambda_{i}^{2}\leq\varepsilon n^{2}.$ Having chosen $\kappa$, set $B=\sum_{\lambda_{i}\geq\kappa n}\lambda_{i}v_{i}v_{i}^{T}$. It satisfies that
$\|A-B\|_{F}^{2}=\sum_{\lambda_{i}<\kappa n}\lambda_{i}^{2}\leq\varepsilon n^{2}.$
Furthermore, the rank of $B$ is at most $\kappa^{-2}$ as $n^{2}\geq\|A\|_{F}^{2}\geq\sum_{\lambda_{i}\geq\kappa n}\lambda_{i}^{2}\geq\operatorname{rank}(B)\cdot(\kappa n)^{2}$. ∎
Given a graph $G$, $\delta\in(0,1)$, and two disjoint sets $X,Y\subset V(G)$, the pair $(X,Y)$ is $\delta$-empty if there are at most $\delta|X||Y|$ edges between $X$ and $Y$. Also, $(X,Y)$ is $\delta$-full if there are at least $(1-\delta)|X||Y|$ edges between $X$ and $Y$. Then $(X,Y)$ is $\delta$-homogeneous if it is either $\delta$-empty or $\delta$-full. A $\delta$-regular partition of $G$ is an equipartition (a partition where all parts share the same size) $V_{1},\dots,V_{K}$ of the vertex set such that all but at most $\delta K^{2}$ of the pairs $(V_{i},V_{j})$ for $1\leq i<j\leq K$ are $\delta$-homogeneous.
@@
For every $\delta\in(0,1)$, there exists $\varepsilon>0$ such that the following holds for every positive integer $r$, and every $n$ that is sufficiently large with respect to $\delta,r$. Let $G$ be a graph with adjacency matrix $A$. Assume that there exists an $n\times n$ symmetric matrix $B$ of rank $r$ such that $\|A-B\|_{F}^{2}\leq\varepsilon n^{2}$. Then $G$ has a $\delta$-regular partition into $K$ parts, where $1/\delta<K<O_{r,\delta}(1)$.
@@
Pick $\beta:=10^{-3}\delta^{1/2}r^{-3/2}$. For $i\in[r]$ and $\ell\in\mathbb{Z}$, let
@@
$\bigcup_{\overline{\ell}\in I}X_{\overline{\ell}}\geq n-r\cdot\frac{\delta n}{8r}\geq n\Big(1-\frac{\delta}{8}\Big).$ (7)
@@
$\eta=\eta_{\overline{\ell}_{1},\overline{\ell}_{2}}=\sum_{i=1}^{r}\mu_{i}\cdot\frac{\beta^{2}}{n}\cdot\overline{\ell}_{1}(i)\overline{\ell}_{2}(i).$
@@
$\left|w_{i}(j_{1})w_{i}(j_{2})-\frac{\beta^{2}}{n}\overline{\ell_{1}}(i)\overline{\ell_{2}}(i)\right|\leq\frac{\beta^{2}}{n}\cdot 4h,$
which we get from the general inequality $|ab-cd|\leq|a||b-d|+|d||a-c|$. Using that $|\mu_{i}|\leq 2n$, we have
$|B_{j_{1},j_{2}}-\eta|\leq\sum_{i=1}^{r}|\mu_{i}|\cdot\left|w_{i}(j_{1})w_{i}(j_{2})-\frac{\beta^{2}}{n}\overline{\ell_{1}}(i)\overline{\ell_{2}}(i)\right|\leq\sum_{i=1}^{r}|\mu_{i}|\cdot\frac{\beta^{2}}{n}\cdot 4h\leq 8r\beta^{2}h<\frac{1}{3}.$
Furthermore, observe that if $X\subset X_{\overline{\ell}_{1}}$ and $Y\subset X_{\overline{\ell}_{2}}$ are such that $(X,Y)$ is not $\delta$-homogeneous, then
$\big\|A[X\times Y]-B[X\times Y]\big\|_{F}^{2}\geq\frac{\delta}{36}|X||Y|.$
Indeed, if $\eta_{\overline{\ell}_{1},\overline{\ell}_{2}}<1/2$, then $A_{j_{1},j_{2}}-B_{j_{1},j_{2}}\geq 1/6$ for every $A_{j_{1},j_{2}}=1$, otherwise $|A_{j_{1},j_{2}}-B_{j_{1},j_{2}}|\geq 1/6$ for every $A_{j_{1},j_{2}}=0$.
Now let $K=|I|/(8\delta)$, and define an equipartition $V_{1},\dots,V_{K}$ of $V(G)$ as follows. To avoid certain technicalities coming from rounding, we assume that $K$ divides $n$. Let $V^{*}$ be the set of elements not covered by any of the $X_{\overline{\ell}}$ for $\overline{\ell}\in I$. For each $\overline{\ell}\in I$, partition $X_{\overline{\ell}}$ arbitrarily into sets, each of size $n/K$, with at most one exceptional set whose size is less than $n/K$. Move the elements of the exceptional set to $V^{*}$. Then finally partition $V^{*}$ into sets of size $n/K$. Let $V_{1},\dots,V_{K}$ be the collection of all the sets in these partitions. Each $X_{\overline{\ell}}$ contributes at most $n/K$ elements to $V^{*}$, so in the end we have $|V^{*}|\leq\delta n/8+|I|\cdot(n/K)\leq\delta n/4$. Therefore, at most $\delta K/4$ sets $V_{i}$ are contained in $V^{*}$. We show that $V_{1},\dots,V_{K}$ is a $\delta$-regular partition.
Assume that $(V_{i},V_{j})$ is not $\delta$-homogeneous. There are at most $\delta K^{2}/2$ such pairs where either $V_{i}\subset V^{*}$ or $V_{j}\subset V^{*}$. On the other hand, if $V_{i},V_{j}\not\in V^{*}$, then $\|A[V_{i}\times V_{j}]-B[V_{i}\times V_{j}]\|_{F}^{2}\geq\frac{\delta}{36}|V_{i}||V_{j}|$. As $\|A-B\|_{F}^{2}\leq\varepsilon n^{2}$, this means that the number of such pairs is at most $36\varepsilon/\delta K^{2}\leq\delta K^{2}/2$. Hence, the total number of pairs that are not $\delta$-homogeneous is at most $\delta K^{2}$, as desired. ∎
An important feature of Lemma 8.2 that $\varepsilon$ only depends on $\delta$, and not on $r$. To continue from this point, we observe that if $X,Y,Z$ are sets of linear sizes such that $(X,Y)$ and $(Y,Z)$ are $\delta$-full, then $(X,Z)$ cannot be $\delta$-empty, assuming $\operatorname{surp}(G)$ is small.